# The Scaling Law in Stellar Light Curves

### v1

@@ 5 Discussion and Conclusion
We demonstrate that with stellar light curves, there is also an emergence of a scaling law, where the GPT-2-like models continue to improve the next-token prediction huber loss with a scaling law of $\log_{10}L_{\text{pretraining}}(C)=-0.00448\log_{10}C-2.386$. Such a scaling law also translates into a more robust and higher level of abstraction in the learned latent representation. The robust latent representation leads to a scaling law on the downstream inference MSE of the surface gravity of stars, with a scaling law of $\log_{10}L_{\text{downstream}}(N)=-0.063\log_{10}\log N-0.725$. With a labeled training set of 10,000 samples and a model size of $>85$ M parameters, the autoregressive models outperform the supervised Transformer model trained from scratch on 100,000 samples, and these results continue to extrapolate to a higher precision with up to $10^{9}$ parameters. At a fixed precision for $\log g$, GPT-2 XL with $1.5$B parameters generative model can achieve such precision consistently with 3-10 times fewer labeled samples compared to the mere supervised learning approach.
@@
Nonetheless, the study demonstrates that the Transformer-based models that have shown wild success in real-world applications also translate into astronomical time series data. The scalability of the models up to 1.5B parameters with our small training set vividly demonstrates the possibility of having a robust and generalizable foundational model for astronomical time series data with sufficient compute. Beyond asteroseismology, the same method can clearly also translate into other forms of astronomical time series data, such as understanding solar flares (Wei et al., 2006) or classifying and finding out-of-distribution transient events (Ivezić et al., 2019).
Ongoing and future surveys, TESS, ZTF, Rubin and SiTian, are estimated to collect observations at scales of 1T, 1P, 14P, and 100P, respectively. There will be a myriad of time series data, perhaps rivaling the text data on the internet and perhaps beyond. The scaling law motivates the dedication of compute resources to characterize astronomical time series data with these models. We forecast that the $\log g$ prediction MSE can be improved two folds than our GPT-2 XL model given $10^{4}$ PF-days, i.e., $\sim 10^{5}$ A100 GPU days. Looking forward, just as with much of the highly non-linear and complex data in real-life scenarios, while the still somewhat simple physical drivers of the light curves of celestial objects can be understood, to a certain extent, with analytical derivations and methods, robust inference and ultimate characterization might still require massive compute. The scalable model demonstrated here is a key step toward unlocking the full potential of astronomical time series data and paving the way for expediting discoveries in the ever-changing dark sky.

### v2

@@ 5 Discussion and Conclusion
We demonstrate that with stellar light curves, there is also an emergence of a scaling law, where the GPT-2-like models continue to improve the next-token prediction huber loss with a scaling law of $\log_{10}L_{\text{pretraining}}(C)=-0.00448\log_{10}C-2.386$. Such a scaling law also translates into a more robust and higher level of abstraction in the learned latent representation. The robust latent representation leads to a scaling law on the downstream inference MSE of the surface gravity of stars, with a scaling law of $\log_{10}L_{\text{downstream}}(N)=-0.063\log_{10}N-0.725$. With a labeled training set of 10,000 samples and a model size of $>85$ M parameters, the autoregressive models outperform the supervised Transformer model trained from scratch on 100,000 samples, and these results continue to extrapolate to a higher precision with up to $10^{9}$ parameters. At a fixed precision for $\log g$ inference, GPT-2 XL with $1.5$B parameters generative model can achieve such precision consistently with 3-10 times fewer labeled samples compared to the mere supervised learning approach.
@@
Nonetheless, the study demonstrates that the Transformer-based models that have shown wild success in real-world applications are also promising for astronomical time series data. The scalability of the models up to 1.5B parameters with our small training set vividly demonstrates the possibility of having a robust and generalizable foundational model for astronomical time series data with sufficient compute. Beyond asteroseismology, the same method can clearly also translate into other forms of astronomical time series data, such as understanding solar flares (Wei et al., 2006) or classifying and finding out-of-distribution transient events (Ivezić et al., 2019).
Ongoing and future surveys, TESS, ZTF, Rubin and SiTian, are estimated to collect observations at scales of 1T, 14T, 14P, and 100P, respectively. There will be a myriad of time series data, perhaps rivaling the text data on the internet and perhaps beyond. The scaling law motivates the dedication of compute resources to characterize astronomical time series data with these models. We forecast that the $\log g$ prediction MSE can be improved two folds than our GPT-2 XL model given $10^{4}$ PF-days, i.e., $\sim 10^{5}$ A100 GPU days. Looking forward, just as with much of the highly non-linear and complex data in real-life scenarios, while the still somewhat simple physical drivers of the light curves of celestial objects can be understood, to a certain extent, with analytical derivations and methods, robust inference and ultimate characterization might still require massive compute. The scalable model demonstrated here is a key step toward unlocking the full potential of astronomical time series data and paving the way for expediting discoveries in the ever-changing dark sky.